# Generative Adversarial Networks for High-Dimensional Item Factor Analysis: A Deep Adversarial Learning Algorithm

### v2

@@ 1 Introduction
An alternative representation paradigm, Generative Adversarial Networks (GANs, Goodfellow et al.,, 2014), a celebrated framework for estimating generative models in deep learning and computer vision, offer greater flexibility in modeling arbitrary probability densities and can contribute to the expressiveness. Multiple variants of GANs have achieved remarkable success in complex tasks such as image generation (e.g., Radford,, 2015; Arjovsky et al.,, 2017; Zhu et al., 2017a, ). Among these models, the combination of GANs and VAEs can be traced back to Adversarial Autoencoders (AAEs; Makhzani et al.,, 2015) and the VAE-GAN model (Larsen et al.,, 2016). Figure 1 provides an intuitive visual example demonstrating how integrating GANs and VAEs can lead to better representation learning. The figure compares ten handwritten images sampled from four models: (a) GAN, (b) WGAN (a GAN variant, Arjovsky et al.,, 2017), (c) VAE, and (d) VAE-GAN. While GAN generated only the digit 1, WGAN exhibited unstable performance, and the images produced by VAE appeared blurry. In contrast, VAE-GAN generated relatively clear results. However, vanilla GANs, VAE-GANs, and AAEs rely solely on adversarial loss, formulated as a minimax game between networks to deceive a discriminator. As a result, they are not motivated from a maximum-likelihood perspective and are therefore not directly applicable for improving Marginal Maximum Likelihood (MML) estimation in IFA.
@@
Therefore, this paper focuses on an improved approach based on VAEs, Adversarial Variational Bayes (AVB, Mescheder et al.,, 2017), which has a flexible and general approach to construct inference models. AVB improves the traditional VAE framework by integrating elements from GANs, specifically by introducing an auxiliary discriminator network. This discriminator engages in an adversarial training process with the encoder to help the inference model produce more accurate and expressive latent variable estimates. By combining the strengths of both VAEs and GANs, AVB removes the restrictive assumption of standard normal distributions in the inference model and enables effective exploration of complex, multimodal latent spaces. Theoretical and empirical evidence has shown that AVB can achieve higher log-likelihood, lower reconstruction error and more precise approximation of the posterior (Mescheder et al.,, 2017). To be specific, trained on a simple synthetic dataset containing samples from 4 different labels, Figure LABEL:fig:toy_comp shows that the posterior for VAE witnessed some empty space between different regions but AVB achieved a complex posterior distribution. This approach not only improves the flexibility and accuracy of MML estimation but also provides more potential for better handling more diverse and high-dimensional data that become increasingly common in psychological and educational research.
@@
The remaining parts are organized as follows. Section 2 explains how AVB builds upon traditional VAE methods for parameter estimation for the Graded Response Model—a general framework within Item Factor Analysis. Section 3 details the implementation steps of the AVB algorithm and introduces strategies for improving its overall performance. Section 4 then provides a comparative evaluation of AVB and VAE, using both empirical and simulated data. Finally, Section 5 examines potential limitations of the proposed approach and highlights directions for future development.
@@ 2.2 Marginal Maximum Likelihood
Marginal Maximum Likelihood estimates model parameters by maximizing the likelihood of the observed data while integrating over latent variables . As a classical approach for parameter estimation in item response models, MML typically employs the Expectation-Maximization (EM) algorithm (e.g., Dempster et al.,, 1977; Bock and Aitkin,, 1981; Muraki,, 1992; Rizopoulos,, 2007; Johnson,, 2007) or quadrature-based methods (e.g., Darrell Bock and Lieberman,, 1970; Rabe-Hesketh et al.,, 2005). The basic concepts of MML estimation are outlined as follows.
@@
Considering all the responses $\bm{x}_{i}=(x_{i,j})_{j=1}^{M}$ of the respondent $i$, since each response is conditionally independent given the latent variables, the conditional probability is: for $\bm{\theta}=(\bm{\theta}_{j})_{j=1}^{M}$,
@@
Given the full $N\times M$ responses matrix $\bm{X}=[x_{i,j}]_{i=1,\dots,N,\,j=1\dots,M}$ of all the respondents, the likelihood of the observed variables is:
@@
Since the above equation contains $N$ integrals in high-dimensional latent space $\mathbb{R}^{P}$, directly maximizing the likelihood is computationally intensive. Instead, variational methods that the current study focuses on improving, approximate $\log\mathcal{L}(\bm{\theta}\mid\bm{X})$ by one lower bound and achieve efficient computation.
@@ 2.3 Variational Inference and Variational Autoencoder
Variational methods approximate the intractable posterior $p_{\bm{\theta}}(\bm{z}\mid\bm{x})$ with a simpler inference model $q_{\bm{\phi}}(\bm{z}\mid\bm{x})$, and have gained increasing attention in the context of IFA (e.g., Cho et al.,, 2021; Natesan et al.,, 2016; Jeon et al.,, 2017; Urban and Bauer,, 2021). For distributions $p$ and $q$, Variational Inference aims to minimize the Kullback-Leibler (KL) divergence, which measures the difference between $p$ and $q$:
[…]
with equality if and only if $p=q$ almost everywhere with respect to $q$.
The marginal log-likelihood of each observation $\bm{x}$, $p_{\bm{\theta}}(\bm{x})$, can be further decomposed into two parts:
$\log p_{\bm{\theta}}(\bm{x})$ $=\text{KL}\left[q_{\bm{\phi}}(\bm{z}\mid\bm{x}),p_{\bm{\theta}}(\bm{z}\mid\bm{x})\right]+\mathbb{E}_{q_{\bm{\phi}}(\bm{z}\mid\bm{x})}\left[\log p_{\bm{\theta}}(\bm{z},\bm{x})-\log q_{\bm{\phi}}(\bm{z}\mid\bm{x})\right]$ (7)
$\geq\mathbb{E}_{q_{\bm{\phi}}(\bm{z}\mid\bm{x})}\left[\log p_{\bm{\theta}}(\bm{z},\bm{x})-\log q_{\bm{\phi}}(\bm{z}\mid\bm{x})\right]$
$=\mathbb{E}_{q_{\bm{\phi}}(\bm{z}\mid\bm{x})}\left[\log p_{\bm{\theta}}(\bm{x}\mid\bm{z})\right]-\text{KL}\left[q_{\bm{\phi}}(\bm{z}\mid\bm{x}),p_{\bm{\theta}}(\bm{z})\right]=\text{ELBO}.$ (8)
The KL divergence term is non-negative, so the expectation term in the Equation (7) is the evidence lower bound (ELBO) since for any $\bm{\theta}$, marginal log-likelihood of the observed data must be greater than ELBO. The ELBO can be further decomposed into two terms in the Equation (8). The first term is the negative reconstruction loss, quantifying the difference between the generated data and the observed data. The second term is the KL divergence between $q_{\bm{\phi}}(\bm{z}\mid\bm{x})$ and $p_{\bm{\theta}}(\bm{z})$, regularizing them to stay close.
For each set of item parameters $\bm{\theta}$, if there exists $\bm{\phi}$ such that $q_{\bm{\phi}}(\bm{z}\mid\bm{x})$ exactly match $p_{\bm{\theta}}(\bm{z}\mid\bm{x})$ the KL divergence term will be zero, so the MML estimator $\bm{\theta}^{*}$ will also be the maximizer of ELBO. However, this will be a very ideal case. Generally speaking, for fixed $\bm{\theta}$, if ELBO is increased by taking gradient steps in $\bm{\phi}$ space, the KL divergence will more approach to its minimum. On the other hand, if $\bm{\phi}$ is fixed, increasing ELBO with respect to $\bm{\theta}$ will push the marginal likelihood $p_{\bm{\theta}}(\bm{x})$ higher. Therefore, our objective is
$\max_{\bm{\theta}}\max_{\bm{\phi}}$ $\mathbb{E}_{p_{\bm{D}}(\bm{x})}\left[-\text{KL}\left[q_{\bm{\phi}}(\bm{z}\mid\bm{x}),p_{\bm{\theta}}(\bm{z})\right]+\mathbb{E}_{q_{\bm{\phi}}(\bm{z}\mid\bm{x})}\log p_{\bm{\theta}}(\bm{x}\mid\bm{z})\right]$ (9)
$=$ $\mathbb{E}_{p_{\bm{D}}(\bm{x})}\mathbb{E}_{q_{\bm{\phi}}(\bm{z}\mid\bm{x})}\left(\log p(\bm{z})-\log q_{\bm{\phi}}(\bm{z}\mid\bm{x})+\log p_{\bm{\theta}}(\bm{x}\mid\bm{z})\right).$
@@
where $\bm{\phi}(\bm{x})=(\bm{\mu}(\bm{x}),\bm{\sigma}(\bm{x}))$ denotes the normal distribution’s configuration in the form of vector function of $\bm{x}$ for convenience; $\bm{x}$, $\bm{\mu}(\bm{x})$, and $\bm{\sigma}(\bm{x})$ are vectors with the same dimension. Given the latent representation $\bm{z}$, the decoder is the probabilistic model which reconstructs or generates data via $p_{\bm{\theta}}(\bm{x}\mid\bm{z})$. However, the inference model expressed in the Equation (10) might not be expressive enough to capture the true posterior (Mescheder et al.,, 2017). The KL divergence term in the ELBO requires explicit forms for both $q_{\bm{\phi}}(\bm{z}\mid\bm{x})$ and $p_{\bm{\theta}}(\bm{z})$, which can limit the choice of posterior families. Highly expressive inference models can lead to higher log-likelihood bounds and stronger decoders making full use of the latent space (Kingma et al.,, 2016; Chen et al.,, 2016).
@@ 2.4 Adversarial Variational Bayes
Adversarial Variational Bayes borrows ideas from the Generative Adversarial Network framework, which leads to remarkable results in areas such as image generation (Goodfellow et al.,, 2014) and data augmentation (Zhu et al., 2017b, ). As shown in the Figure 3, GANs employ adversarial training by pitting a generator against a discriminator, compelling each to enhance its performance. This competition results in highly convincing synthetic data, highlighting GANs’ power in discovering intricate patterns found in real-world samples.
Figure 3: Schematic illustration of the general Generative Adversarial Networks.
To integrate GANs into VAE, AVB regards the inference model $q_{\bm{\phi}}(\bm{z}\mid\bm{x})$ as the generator of latent variables and replaces the explicit KL divergence computation in the VAE algorithm with an adversarially trained discriminator, allowing for more expressive inference models. To be specific, besides the Encoder network $q_{\bm{\phi}}(\bm{z}\mid\bm{x})$ and the Decoder network $p_{\bm{\theta}}(\bm{x}\mid\bm{z})$, a neural network $T(\bm{x},\bm{z})$ as the Discriminator learns to distinguish between samples from the inference model $q_{\bm{\phi}}(\bm{z}\mid\bm{x})$ and the prior $p_{\bm{\theta}}(\bm{z})$.
Consider the following objective for the discriminator $T(\bm{x},\bm{z})$ for a given $q_{\bm{\phi}}(\bm{z}\mid\bm{x})$:
$\max_{T}\,\mathbb{E}_{p_{\bm{D}}(\bm{x})}\mathbb{E}_{q_{\bm{\phi}}(\bm{z}\mid\bm{x})}\log\sigma\left(T(\bm{x},\bm{z})\right)+\mathbb{E}_{p_{\bm{D}}(\bm{x})}\mathbb{E}_{p(\bm{z})}\log\left(1-\sigma\left(T(\bm{x},\bm{z})\right)\right).$ (11)
Here, $\sigma(t):=\left(1+e^{-t}\right)^{-1}$ denotes the sigmoid-function. Intuitively, $T(\bm{x},\bm{z})$ tries to distinguish pairs $(\bm{x},\bm{z})$ that are sampled independently using the distribution $p_{\bm{D}}(\bm{x})p(\bm{z})$ from those that are sampled using the current inference model (i.e., $p_{\bm{D}}(\bm{x})q_{\bm{\phi}}(\bm{z}\mid\bm{x})$). The Equation (11) encourages samples from $q_{\bm{\phi}}(\bm{z}\mid\bm{x})$ but punishes samples from $p(\bm{z})$. According to Mescheder et al., (2017), the KL term in the ELBO can be recovered by the following theorem.
Theorem 2.1.
@@
so $\text{KL}\left[q_{\bm{\phi}}(\bm{z}\mid\bm{x}),p_{\bm{\theta}}(\bm{z})\right]=\mathbb{E}_{q_{\bm{\phi}}(\bm{z}\mid\bm{x})}\left(T^{*}(\bm{x},\bm{z})\right)$.
@@
$\max_{\bm{\theta}}\max_{\bm{\phi}}$ $\mathbb{E}_{p_{\bm{D}}(\bm{x})}\mathbb{E}_{q_{\bm{\phi}}(\bm{z}\mid\bm{x})}\left(-T^{*}(\bm{x},\bm{z})+\log p_{\bm{\theta}}(\bm{x}\mid\bm{z})\right)$ (13)
$=$ $\mathbb{E}_{p_{\bm{D}}(\bm{x})}\mathbb{E}_{\bm{\epsilon}}\left(-{T}^{*}\left(\bm{x},\bm{z}_{\bm{\phi}}(\bm{x},\bm{\epsilon})\right)+\log p_{\bm{\theta}}\left(\bm{x}\mid\bm{z}_{\bm{\phi}}(\bm{x},\bm{\epsilon})\right)\right).$
Therefore, we need to do two optimizations for Objectives (11) and (13), summarized in Algorithm 1. Note that $T(\bm{x},\bm{z})$ and $\bm{z}_{\bm{\phi}}(\bm{x},\bm{\epsilon})$ can be implemented as neural networks, but in the context of the GRM, the decoder $p_{\bm{\theta}}\left(\bm{x}\mid\bm{z}_{\bm{\phi}}(\bm{x},\bm{\epsilon})\right)$ corresponds directly to the likelihood definition for the GRM. As such, it is modeled by a single-layer neural network consisting of a linear function and a sigmoid activation.

### v3

@@ 1 Introduction
An alternative representation paradigm, Generative Adversarial Networks (GANs, Goodfellow et al.,, 2014), a celebrated framework for estimating generative models in deep learning and computer vision, offer greater flexibility in modeling arbitrary probability densities and can contribute to the expressiveness. Multiple variants of GANs have achieved remarkable success in complex tasks such as image generation (e.g., Radford,, 2015; Arjovsky et al.,, 2017; Zhu et al., 2017a, ). Among these models, the combination of GANs and VAEs can be traced back to Adversarial Autoencoders (AAEs; Makhzani et al.,, 2015) and the VAE-GAN model (Larsen et al.,, 2016). Figure 1 provides an intuitive visual example demonstrating how integrating GANs and VAEs can lead to better representation learning. The figure compares ten handwritten images sampled from four models: (a) GAN, (b) WGAN (a GAN variant, Arjovsky et al.,, 2017), (c) VAE, and (d) VAE-GAN. While GAN generated only the digit 1111The generator’s output of only ‘1’s indicates mode collapse, a common failure mode in vanilla GAN training. The standard GAN objective encourages the network to repeatedly generate identical samples that incur minimal loss, rather than to explore diverse outputs that might be more heavily penalized (Mi et al.,, 2018)., WGAN exhibited unstable performance, and the images produced by VAE appeared blurry. In contrast, VAE-GAN generated relatively clear results for most of digits. However, vanilla GANs, VAE-GANs, and AAEs rely solely on adversarial loss, formulated as a minimax game between networks to deceive a discriminator. As a result, they are not motivated from a maximum-likelihood perspective and are therefore not directly applicable for improving Marginal Maximum Likelihood (MML) estimation in IFA.
@@
Therefore, this paper focuses on an improved approach based on VAEs, Adversarial Variational Bayes (AVB, Mescheder et al.,, 2017), which has a flexible and general approach to construct inference models. AVB improves the traditional VAE framework by integrating elements from GANs, specifically by introducing an auxiliary discriminator network. This discriminator engages in an adversarial training process with the encoder to help the inference model produce more accurate and expressive latent variable estimates. By combining the strengths of both VAEs and GANs, AVB removes the restrictive assumption of standard normal distributions in the inference model and enables effective exploration of complex, multimodal latent spaces. Theoretical and empirical evidence has shown that AVB can achieve higher log-likelihood, lower reconstruction error and more precise approximation of the posterior (Mescheder et al.,, 2017). For example, we consider a simple dataset containing samples from 4 different labels whose latent codes are uniformly distributed. Figure 3(a) compares the recovered latent distribution: the unallocated regions (“empty space”) and overlap between regions in VAE’s result can yield ambiguous latent codes and compromise both inference precision and the quality of generated samples, but AVB produces a posterior with well-defined boundaries around each region. This approach not only improves the flexibility and accuracy of MML estimation but also provides more potential for better handling more diverse and high-dimensional data that become increasingly common in psychological and educational research.
(a) Distribution of latent variables for VAE and AVB trained on a simple synthetic dataset containing samples from 4 different labels.
@@
The remaining parts are organized as follows. Section 2 explains how AVB builds upon traditional VAE methods for parameter estimation for the Graded Response Model, a general framework within Item Factor Analysis. Section 3 details the implementation steps of the AVB algorithm and introduces strategies for improving its overall performance. Section 4 then provides a comparative evaluation of AVB and VAE, using both empirical and simulated data. Finally, Section 5 examines potential limitations of the proposed approach and highlights directions for future development.
@@ 2.2 Marginal Maximum Likelihood
Marginal Maximum Likelihood estimates model parameters by maximizing the likelihood of the observed response data while integrating over latent variables . As a classical approach for parameter estimation in item response models, MML typically employs the Expectation-Maximization (EM) algorithm (e.g., Dempster et al.,, 1977; Bock and Aitkin,, 1981; Muraki,, 1992; Rizopoulos,, 2007; Johnson,, 2007) or quadrature-based methods (e.g., Darrell Bock and Lieberman,, 1970; Rabe-Hesketh et al.,, 2005). The basic concepts of MML estimation are outlined as follows.
@@
Considering the response pattern $\bm{x}_{i}=(x_{i,j})_{j=1}^{M}$ of the respondent $i$, since each response is conditionally independent given the latent variables, the conditional probability is: for $\bm{\theta}=(\bm{\theta}_{j})_{j=1}^{M}$,
@@
Given the full $N\times M$ responses matrix $\bm{X}=[x_{i,j}]_{i=1,\dots,N,\,j=1\dots,M}$ of all the respondents, the MML estimator $\bm{\theta}^{*}$ of item parameter is achieved by maximizing the likelihood of the observed responses:
@@
Since the above equation contains $N$ integrals in high-dimensional latent space $\mathbb{R}^{P}$, directly maximizing the likelihood is computationally intensive. Instead, variational methods that the current study focuses on improving, approximate $\log\mathcal{L}(\bm{\theta}\mid\bm{X})$ by one lower bound and achieve efficient computation. This lower bound is derived based on the latent variable posterior $p_{\bm{\theta}}(\bm{z}\mid\bm{x})=p_{\bm{\theta}}(\bm{x}\mid\bm{z})p(\bm{z})/p_{\bm{\theta}}(\bm{x})$.
@@ 2.3 Variational Inference and Variational Autoencoder
Variational methods approximate the intractable posterior $p_{\bm{\theta}}(\bm{z}\mid\bm{x})$ with a simpler inference model $q_{\bm{\phi}}(\bm{z}\mid\bm{x})$, and have gained increasing attention in the context of IFA (e.g., Cho et al.,, 2021; Jeon et al.,, 2017; Urban and Bauer,, 2021). For distributions $p$ and $q$, Variational Inference aims to minimize the Kullback-Leibler (KL) divergence, ${\text{KL}}\left(q,p\right)$ [^2] The KL divergence is defined as ${\text{KL}}\left(q,p\right)=\mathbb{E}_{q}\left[\log q-\log p\right]=\int q_{\bm{\phi}}(\bm{z}\mid\bm{x})\left[\log q_{\bm{\phi}}(\bm{z}\mid\bm{x})-\log p_{\bm{\theta}}(\bm{z}\mid\bm{x})\right]d\bm{z}\geq 0,$ with equality if and only if $p=q$ almost everywhere with respect to $q$., which measures the difference between $p$ and $q$. This is related to the marginal log-likelihood of each response pattern $\bm{x}$, due to following decomposition:
$\log p_{\bm{\theta}}(\bm{x})=\text{KL}\left[q_{\bm{\phi}}(\bm{z}\mid\bm{x}),p_{\bm{\theta}}(\bm{z}\mid\bm{x})\right]+\underbrace{\mathbb{E}_{q_{\bm{\phi}}(\bm{z}\mid\bm{x})}\left[\log p_{\bm{\theta}}(\bm{z},\bm{x})-\log q_{\bm{\phi}}(\bm{z}\mid\bm{x})\right]}_{\text{ELBO}(\bm{x})}$ (6)
$\text{KL}(q,p)$ is non-negative, so the expectation term is the evidence lower bound (ELBO). For fixed $\bm{\theta}$, maximizing ELBO in $\bm{\phi}$ space is equivalent to minimizing $\text{KL}(q,p)$, while for fixed $\bm{\phi}$, increasing ELBO with respect to $\bm{\theta}$ will push the marginal likelihood $p_{\bm{\theta}}(\bm{x})$ higher. The best case is that if $q_{\bm{\phi}}(\bm{z}\mid\bm{x})$ exactly matches $p_{\bm{\theta}}(\bm{z}\mid\bm{x})$, $\text{KL}(q,p)=0$ and the MML estimator $\bm{\theta}^{*}$ will also be the maximizer of ELBO. The ELBO can be further decomposed:
$\text{ELBO}(\bm{x})=\mathbb{E}_{q_{\bm{\phi}}(\bm{z}\mid\bm{x})}\left[\log p_{\bm{\theta}}(\bm{x}\mid\bm{z})\right]-\text{KL}\left[q_{\bm{\phi}}(\bm{z}\mid\bm{x}),p(\bm{z})\right]$ (7)
The first term is the negative reconstruction loss, quantifying how accurately the estimated GRM can get the original input response $\bm{x}$ from the respondent’s latent $\bm{z}$. The second term is the KL divergence between $q_{\bm{\phi}}(\bm{z}\mid\bm{x})$ and $p(\bm{z})$, regularizing them to stay close. Considering the empirical distribution of the observed item response patterns, $p_{\bm{D}}(\bm{x})$, our objective, $max_{\bm{\theta},\bm{\phi}}\;\mathbb{E}_{p_{\bm{D}}(\bm{x})}\left[\text{ELBO}(\bm{x})\right]$, can be further derived as:
$\max_{\bm{\theta}}\max_{\bm{\phi}}\;\mathbb{E}_{p_{\bm{D}}(\bm{x})}\mathbb{E}_{q_{\bm{\phi}}(\bm{z}\mid\bm{x})}\left(\log p(\bm{z})-\log q_{\bm{\phi}}(\bm{z}\mid\bm{x})+\log p_{\bm{\theta}}(\bm{x}\mid\bm{z})\right).$ (8)
@@
where $\bm{\phi}(\bm{x})=(\bm{\mu}(\bm{x}),\bm{\sigma}(\bm{x}))$ denotes the normal distribution’s configuration in the form of vector function of $\bm{x}$ for convenience; $\bm{x}$, $\bm{\mu}(\bm{x})$, and $\bm{\sigma}(\bm{x})$ are vectors with the same dimension. Given the latent representation $\bm{z}$, the decoder is the probabilistic model which reconstructs or generates data via $p_{\bm{\theta}}(\bm{x}\mid\bm{z})$. However, the inference model expressed in the Equation (9) might not be expressive enough to capture the true posterior (Mescheder et al.,, 2017). The KL divergence term in the ELBO requires explicit forms for both $q_{\bm{\phi}}(\bm{z}\mid\bm{x})$ and $p(\bm{z})$, which can limit the choice of posterior families. Highly expressive inference models can lead to higher log-likelihood bounds and stronger decoders making full use of the latent space (Kingma et al.,, 2016; Chen et al.,, 2016). One remedy is to replace the inference model with an implicit distribution that lacks a closed-form density and to approximate the intractable $\text{KL}\left[q_{\bm{\phi}}(\bm{z}\mid\bm{x}),p(\bm{z})\right]$ via auxiliary networks (Shi et al.,, 2018). AVB is a representative example of this approach.
@@ 2.4 Adversarial Variational Bayes
Adversarial Variational Bayes borrows ideas from the Generative Adversarial Network framework, which leads to remarkable results in areas such as image generation (Goodfellow et al.,, 2014) and data augmentation (Zhu et al., 2017b, ). As shown in the Figure 3, GANs employ adversarial training by pitting a generator against a discriminator, compelling each to enhance its performance. This competition results in highly convincing synthetic data, highlighting GANs’ power in discovering intricate patterns found in real-world samples. In the context of IFA, item response data are inputs to both the generator and discriminator. Random noise sampled from the prior distribution is also required for the generator to get “generated” latent variables, which, along with “true” samples drawn directly from the prior distribution, are then evaluated by the discriminator.
Figure 3: Schematic illustration of a standard Generative Adversarial Network. In some GAN variants, real data serve only as true samples and are not fed into the generator. However, in the AVB application to IFA, the generator and discriminator take item response data as input, and the discriminator distinguishes between samples in the latent space.
To integrate GANs into VAE, AVB regards the inference model $q_{\bm{\phi}}(\bm{z}\mid\bm{x})$ as the generator of latent variables and replaces the explicit KL divergence computation in the VAE with an adversarially trained discriminator, allowing for more expressive inference models. To be specific, besides the Encoder network $q_{\bm{\phi}}(\bm{z}\mid\bm{x})$ and the Decoder network $p_{\bm{\theta}}(\bm{x}\mid\bm{z})$, a neural network $T_{\bm{\psi}}(\bm{x},\bm{z})$ with parameters $\bm{\psi}$ as the Discriminator learns to distinguish between samples from the inference model $q_{\bm{\phi}}(\bm{z}\mid\bm{x})$ and the prior $p(\bm{z})$ via minimizing the binary classification loss, which can be formulated as the following objective:
$\max_{\bm{\psi}}\,\mathbb{E}_{p_{\bm{D}}(\bm{x})}\mathbb{E}_{q_{\bm{\phi}}(\bm{z}\mid\bm{x})}\log\sigma\left(T_{\bm{\psi}}(\bm{x},\bm{z})\right)+\mathbb{E}_{p_{\bm{D}}(\bm{x})}\mathbb{E}_{p(\bm{z})}\log\left(1-\sigma\left(T_{\bm{\psi}}(\bm{x},\bm{z})\right)\right).$ (10)
Here, $\sigma(t):=\left(1+e^{-t}\right)^{-1}$ denotes the sigmoid-function. Intuitively, $T_{\bm{\psi}}(\bm{x},\bm{z})$ tries to distinguish pairs $(\bm{x},\bm{z})$ that are sampled independently using the distribution $p_{\bm{D}}(\bm{x})p(\bm{z})$ from those that are sampled using the current inference model (i.e., $p_{\bm{D}}(\bm{x})q_{\bm{\phi}}(\bm{z}\mid\bm{x})$). The Equation (10) encourages samples from $q_{\bm{\phi}}(\bm{z}\mid\bm{x})$ but punishes samples from $p(\bm{z})$. As shown by Mescheder et al., (2017), the KL term in the ELBO can be recovered by the following theorem.
Theorem 2.1 (Mescheder et al.,, 2017).
@@
so $\text{KL}\left[q_{\bm{\phi}}(\bm{z}\mid\bm{x}),p(\bm{z})\right]=\mathbb{E}_{q_{\bm{\phi}}(\bm{z}\mid\bm{x})}\left(T^{*}(\bm{x},\bm{z})\right)$.
@@
$\max_{\bm{\theta}}\max_{\bm{\phi}}\;\mathbb{E}_{p_{\bm{D}}(\bm{x})}\mathbb{E}_{\bm{\epsilon}}\left(-{T}^{*}\left(\bm{x},\bm{z}_{\bm{\phi}}(\bm{x},\bm{\epsilon})\right)+\log p_{\bm{\theta}}\left(\bm{x}\mid\bm{z}_{\bm{\phi}}(\bm{x},\bm{\epsilon})\right)\right)$ (12)
Therefore, we need to do two optimizations for Objectives (10) and (12), summarized in Algorithm 1. Note that $T_{\bm{\psi}}(\bm{x},\bm{z})$ and $\bm{z}_{\bm{\phi}}(\bm{x},\bm{\epsilon})$ can be implemented as neural networks, but in the context of the GRM, the decoder $p_{\bm{\theta}}\left(\bm{x}\mid\bm{z}_{\bm{\phi}}(\bm{x},\bm{\epsilon})\right)$ corresponds directly to the likelihood definition for the GRM. As such, it is modeled by a single-layer neural network consisting of a linear function and a sigmoid activation.